OpenStudy (anonymous):

PLEASE HELP WILL GIVE A MEDAL Find the dot product, a • b. a = 6i + 5j, b = -5i + 4j

OpenStudy (anonymous):

50 <1, 9> -10 <-30, 20>

OpenStudy (jdoe0001):

\(\large \bf <a,b>\cdot <c,d>\implies (a\cdot c)+(b\cdot d)\)

OpenStudy (anonymous):

@jdoe0001 but they dont give it in ordered pair form?